# A Set of Rules for Model Validation

### v1

@@ 1 Introduction
Model validation is a fundamental task in all modern data-driven systems, whether they fall under the broad categories of Statistics, Machine Learning (ML), Artificial Intelligence (AI), or more specialized fields like chemometrics. Validation has become a major focus for regulatory and standardization bodies, with key reports and standards highlighting the growing concern for ensuring the trustworthiness and reliability of data-driven models:
@@
    NIST AI Risk Management Framework (AI RMF 1.0, 2023): Published by the U.S. Department of Commerce, this framework provides management techniques to address the risks and ensure the trustworthiness of AI systems, with validation as a core component.
@@
    The IEEE P2841™-2022 is a recommended practice for the framework and process for deep learning evaluation.
@@
All previous literature is either regulatory or methodological, focusing on a list of specific data processing methods and good practices, like the use of external validation and/or cross-validation (CV). The approach of this perspective paper is different: to suggest a reduced number of conceptual, general rules for proper validation, which provide the intuition behind the methods that are discussed elsewhere. These rules are useful to:
@@
    Design sound validation schemes in complex problems/data pipelines.
@@
    Report transparent and comparable performance results.
@@
    Specify limitations of reported validation results, in terms of different forms of risks.
@@ 2 Rules of Validation
We need to start with the definition of the term validation. In the field of classical (frequentist) statistics, validation is often interpreted as the evaluation of the stability and statistical significance of the model parameters [7]. Thus, the model is valid if the parameters are statistically significant (and so, unlikely to be the result of a purely random process) and stable (they would not change dramatically in front of expected changes in the input). This paper, however, focuses on the more popular definition of validation in the field of machine learning, interpreted as the evaluation of the model’s capability for performance generalization. A valid model is the one that maintains a high modeling performance when applied to new (independent) data from the population of interest, that is, the (potentially unlimited) set of data objects for which the model makes some form of inference (prediction, classification, hypothesis test, anomaly detection, etc.) This second definition also reflects the relevance of stability (when applied to new data) but substitutes statistical significance with performance generalization. Note that these three interconnected concepts: stability, statistical significance and performance generalization, have a principal role in the proposed validation rules.
@@
Data-driven models have two types of parameters: the regular ones, which values are automatically derived from data using a fitting algorithm, and the meta-parameters, which values are selected by the analyst in a pseudo-automatic fashion. For instance, in Principal Component Analysis (PCA) we may use the NIPALS algorithm to fit loadings and scores (the regular parameters), and CV to select the number of components (the single PCA meta-parameter) [8, 9]. Thus, model building comprises two interrelated tasks:
@@
    Training or data fitting. Both terms are interchangeably employed by the ML community, while in chemometrics the term ’calibration’ is popular due to the wide interest in extracting quantitative information from complex analytical instruments, which have to be calibrated. Training, data fitting or calibration refer to the estimation of regular parameters using a suitable fitting algorithm.
@@
    Model selection. Select the model with the best performance among a set of possibilities, which vary in the meta-parameters’ values.
The whole process of model building can be done with two data splits (often called ‘training’ and ‘validation’ sets) or through some form of CV. Once the model is built, the generalization performance is then evaluated with an additional ‘test’ set or through double/nested CV [10]. The separation between model building and test data is a cornerstone of the engineering of data-driven models, since there is often a significant gap between performances in validation and test sets [2], so that performance is often higher in the former. When such a gap exists, we say that the model is overfit to the model building data, capturing patterns that are only found in these data but do not generalize to the test set and to the population of interest.
@@
As Esbense and Geladi claim [12, 13] "All prediction models must be validated with respect to realistic future circumstances". This statement is particularly relevant in terms of the completeness of the test data and its level of independence with the model building data, among other considerations that affect data quality [14].
@@
An important element of a validation strategy that mimics the real-life application is that any relevant data processing operation during model building (either at model fitting or model selection) should not consider the ’test’ data in any way possible. This is simply because in real-life practice, the data that will be inputted into the model is not available in advance, and test data should have the same consideration of no availability. Processing operations at model building which should be blind to the test set often include preprocessing options like mean-centering or auto-scaling (where means and scales can be considered model parameters), variable selection (selected variables can be considered meta-parameters), and similar operations [4]. This also means that nested CV routines should be designed to perform these operations within the inner loop, without using the corresponding ’test’ split of data. An unfortunately extended bad practice [4, 6] is applying variable selection using the entire dataset and subsequently split the data in model building and test. An equivalent error is to perform variable selection and subsequently perform inferential procedures (like t-tests, ANOVA tests, univariate estimates of precision or recall, ROC curves, etc.) on the selected variables, since these procedures do not consider that the variables have been pre-selected from a larger group as the most promising ones.
@@
Quite often, the data used in a study is provided as a single block and we resort to data splitting [4] for the generation of a model building dataset (potentially divided in training and validation) and a test dataset. Single and nested CV is a form of data splitting. Please, note that in this situation, the test data is only independent to the model building data in terms of all the data processing operations that we make AFTER data splitting, but it is not independent in terms of all the previous operations, included the ones involved in the data generation process. This is a limitation that should be recognized.
@@
There is a large number of criteria for performance evaluation [17], to give some few examples: Predictive Error Sum-of-Squares (PRESS), Q2, Mean Absolute Error (MAE), Mean Square Error (MSE), Precision, Recall, F1 Score, Area Under the Receiver Operating Characteristic (AUROC), Number of Misclassification (NMC), Mathews Coefficient Correlation (MCC), Cohen’s Kappa, etc. Quite often, when comparing multiple model variants, these criteria provide conflicting results. In real-life applications, however, there are considerations to make, often of practical nature, about when a solution is preferred over other solutions. For instance, there may be different perceived degrees of severity when making false positives and false negatives. In the diagnosis of a severe disease, a false positive may be an unpleasant experience for the patient, but a false negative may lead to very negative consequences for their health. In a critical infrastructure, a false positive of malfunctioning may lead to wasting the time of a technician, but a false negative may lead to the disruption of the critical service. Even two different types of false positives (or negatives) may not be perceived the same in a real application. Understanding the severity of inference errors in the real-life application of a model is extremely important for proper validation. The application experts should provide this information.
@@
Take the example of Figure 1, where a random $\mathbf{y}~(1000\times 1)$ is estimated by a hypothetical classifier $\hat{\mathbf{y}}~(1000\times 1)$ with some level of noise. We repeat the generation of $\mathbf{y}$ and $\hat{\mathbf{y}}$ 10 times, emulating what would happen if we assess the performance of the classifier over 10 different test datasets with the same imbalance level. When $\mathbf{y}$ is composed of a majority class with $70\%$ of the observations and a minority class with $30\%$, the resulting ROC curve is very stable (Figure 1(a), with an AUROC [^2] The Area Under the ROC curve or AUROC ranges from 0.5 for a random classifier to an optimal value of 1 for a perfect classifier. between 0.84 and 0.86). When we assess the same classifier with a minority class of $1\%$, which is realistic or even large for many anomaly detection problems or in the diagnosis of many diseases, the random variability associated to the ROC curve grows severely (Figure 1(b), with an AUROC between 0.77 and 0.89). Clearly, a ROC (or AUROC) is an unstable performance measure for this type of problems.
@@
Figure 1: Receiver Operating Characteristic (ROC) curves for unbalanced data.
@@
A different angle is considered in Figure 3. Rather than the NMC, we consider a weighted performance function where false positives are pondered with 1 unit and false negatives with 100 units, incorporating the idea that a false negative is something we would rather avoid at the expense of making more false positives. Back in the case with a minority class of $1\%$, this approach shows a better performance of the original classifier (in Figure 3(a)) than the naive approach that always outputs the negative class (in Figure 3(b)). This perceived performance is more consistent for a real-case problem where false negative are more relevant. Furthermore, this performance function shows that reducing the threshold for the detection (which generates less false negatives at the expense of generating more false positives) can further improve performance (Figure 3(c)). Clearly, the nature of the performance function can change both our perception of performance and the consequent choices we make to configure the optimal model. For this reason, choosing a performance criterion that is consistent with the real-life application is key to make adequate choices and have realistic assessment of the model quality.
@@ 2.4 Rule 4: Relevant baselines should always be provided
It is also a strong recommendation to report relevant performance baselines. A baseline can be selected from different perspectives. To give an example, popular CV curves for the selection of the number of latent variables (LVs) in component models (like PCA or Partial Least Squares (PLS)) are too often shown without a baseline. In this case, a suggested baseline [18] can be the cross-validated performance criterion (e.g., the sum of squares) of the model with 0 LVs, that is, only considering the preprocessing parameters, like for instance the cross-validated means. Quite often, not a single model variant, regardless the number of LVs, yields an improvement over the 0 LVs baseline. This is a situation in which the selected component model is not an adequate choice, or there is simply no pattern to be found in the data. This situation cannot be detected if a baseline is not provided.
Take the example of Figure 4, with the CV profile of a PLS model where $\mathbf{X}~(20\times 10)$ and $\mathbf{y}~(20\times 1)$ are unrelated. Figure 4(a) shows a CV curve with a minimum at 2 LVs. An obvious baseline we are missing in this first plot is to fix the minimum of the Y scale at PRESS = 0. When we add it in Figure 4(b) we realize that the difference in PRESS between all the models, especially the ones with 1 LV and 2 LVs, is very small. Adding the 0 LVs baseline, represented by the error of estimating $\mathbf{y}$ with its own (cross-validated) average, also provides useful information. This is illustrated in Figure 4(c), where we also show with dashed lines two hypothetical values for the PRESS at 0 LVs that would dramatically change our interpretation. If the PRESS at 0 LV is 100, the interpretation would be that the model with 1 or 2 LVs is actually capturing more than 60% of the variance in $\mathbf{y}$ in a predictive way. If the PRESS at 0 LV is 10, the interpretation would be that no PLS model has any predictive power. This is actually what the real value of the PRESS at 0 LVs (around 30) is implying.
@@
Another form of baseline is represented by null examples. A null example is a negative case, fabricated using random number generators while maintaining as much as possible the characteristics of the real data at hand. The null distribution in permutation testing [19], or the null effect size in powercurves [20], for instance, follow this intuition. The resulting performance reflects a baseline, which is often (much) higher than expected [21]. I strongly recommend assessing any data pipeline with one or several null examples as an excellent supplement for validation. Null examples are a perfect tool to detect data leakage problems, because they lead to an unexpectedly high performance when such problems exist.
Take now the example of Figure 5, similar as the previous example but for $\mathbf{X}~(20\times 1000)$ and $\mathbf{y}~(20\times 1)$. Since both blocks of data are unrelated, this represent a null example of a case where $\mathbf{X}$ is highly dimensional. A correct validation scheme should reflect the negative prediction capability of the null $\mathbf{X}$. Figure 5(a) shows the (incorrect) case where we do variable selection with the whole dataset and then perform double CV. Figure 5(b) shows the (correct) case where we do variable selection within the inner loop of the double CV. In both cases, we present one example of the CV curve in the inner loop, and the final $Q^{2}$ of the selected model. The null example shows the overly optimistic nature of the incorrect validation scheme in Figure 5(a), where the optimal model has 1 LV and a fairly good (perceived) prediction performance. The correct validation scheme in Figure 5(b) shows that $\mathbf{X}$ does not contain information on $\mathbf{y}$. The nice property of a null example is that it is extremely easy to generate (either from random data or by permutation of a real dataset) but very powerful to detect flawed validation schemes. In very complex pipelines, with several steps for preprocessing, feature engineering and modeling, running some permuted examples is an effortless way to double-check that the pipeline is not suffering from any data leakage problem.
@@
Figure 5: Cross-validation curve and $Q^{2}$ of double cross-validation for Partial Least Squares (PLS) in a simulated dataset where $\mathbf{X}(20\times 1000)$ and $\mathbf{y}(20\times 1)$ are unrelated: variable selection performed before validation (a) and variable selection performed within the inner loop (b).

### v2

@@ 1 Introduction
Model validation is a fundamental task in all modern data-driven systems, whether they fall under the broad categories of Statistics, Machine Learning, Artificial Intelligence, or more specialized fields like chemometrics. Validation has become a major focus for regulatory and standardization bodies, with key reports and standards highlighting the growing concern for ensuring the trustworthiness and reliability of data-driven models:
@@
    NIST Artificial Intelligence Risk Management Framework (AI RMF 1.0, 2023): Published by the U.S. Department of Commerce, this framework provides management techniques to address the risks and ensure the trustworthiness of AI systems, with validation as a core component.
@@
    The American Society for Testing and Materials (ASTM) E2617 provides a formal framework for the validation of empirically derived multivariate calibrations/models.
@@
All previous literature is either regulatory or methodological, focusing on a list of specific data processing methods and good practices, like the use of external validation and/or cross-validation. The approach of this perspective paper is different: to suggest a reduced number of conceptual, general rules for proper validation. These rules provide the underlying intuition for the methods discussed elsewhere. Consequently, they can be leveraged to understand when and why these methods are relevant, or to derive new validation techniques when they are not. In particular, the rules are meant to be useful for:
@@
    Designing sound validation schemes in complex problems/data pipelines [^2] In this paper, ”data pipeline” refers to the set of processing steps performed on the data for its analysis and modeling (e.g., data cleaning, preprocessing, analysis, modeling).
@@
    Reporting transparent and comparable performance results.
@@
    Specifying limitations of reported validation results, in terms of different forms of risks.
@@ 2 Rules of Validation
We need to start with the definition of the term validation. In the field of classical statistics, validation is often interpreted as the evaluation of the stability and statistical significance of the model parameters [7]. Thus, the model is valid if the parameters are statistically significant—and therefore unlikely to be the result of a purely random process—and stable, meaning they would not change dramatically in response to expected variations in the input. This paper, however, focuses on the more popular definition of validation in the field of machine learning, interpreted as the evaluation of the model’s capability for performance generalization. A valid model is the one that maintains a high modeling performance when applied to new independent data from the population of interest, that is, the potentially unlimited set of data objects for which the model makes some form of inference—prediction, classification, hypothesis test, anomaly detection, etc. This second definition also reflects the relevance of stability when applied to new data, but substitutes statistical significance with performance generalization. Note that these three interconnected concepts: stability, statistical significance and performance generalization, have a principal role in the proposed validation rules.
@@
Data-driven models have two types of parameters: the regular ones, which values are automatically derived from data using a fitting algorithm, and the meta-parameters, which values are selected by the analyst in a pseudo-automatic fashion. For instance, in Principal Component Analysis (PCA) we may use the NIPALS algorithm to fit loadings and scores (the regular parameters), and cross-validation to select the number of components (the single PCA meta-parameter) [8, 9]. Thus, model building comprises two interrelated tasks:
@@
    Training or data fitting. Both terms are interchangeably employed by the Machine Learning community, while in chemometrics the term ’calibration’ is popular due to the wide interest in extracting quantitative information from complex analytical instruments, which have to be calibrated. Training, data fitting or calibration refer to the estimation of regular parameters using a suitable fitting algorithm.
@@
    Model selection. From the set of model variants with different meta-parameters, select the model with the best performance.
The whole process of model building can be done with two data splits (often called ‘training’ and ‘validation’ sets) or through some form of cross-validation. Once the model is built, the generalization performance is then evaluated with an additional ‘test’ set or through double/nested cross-validation [10]. The separation between model building and test data is a cornerstone of the engineering of data-driven models, since there is often a significant gap between performances in validation and test sets [2], so that performance is often higher in the former. When such a gap exists, we say that the model is overfit to the model building data, capturing patterns that are only found in these data but do not generalize to the test set and to the population of interest.
@@
As Esbensen and Geladi claim [12, 13] "All prediction models must be validated with respect to realistic future circumstances". This statement is particularly relevant in terms of the completeness of the test data and its level of independence with the model building data, among other considerations that affect data quality [14].
@@
An important element of a validation strategy that mimics the real-life application is that any relevant data processing operation during model building (either at model fitting or model selection) should not consider the ’test’ data in any way possible. This is simply because in real-life practice, the data that will be inputted into the model is not available in advance, and test data should have the same consideration of no availability. Processing operations at model building which should be blind to the test set often include preprocessing options like mean-centering or auto-scaling (where means and scales can be considered model parameters), variable selection (selected variables can be considered meta-parameters), and similar operations [4]. This also means that nested cross-validation routines should be designed to perform these operations within the inner loop, without using the corresponding ’test’ split of data. An unfortunately extended bad practice [4, 6] is applying variable selection using the entire dataset and subsequently split the data in model building and test. An equivalent error is to perform variable selection and subsequently perform inferential procedures (like t-tests, ANOVA tests, univariate estimates of precision or recall, ROC curves, etc.) on the selected variables, since these procedures do not consider that the variables have been pre-selected from a larger group as the most promising ones.
@@
Quite often, the data used in a study is provided as a single block and we resort to data splitting [4] for the generation of a model building dataset (potentially divided in training and validation) and a test dataset. Single and nested cross-validation is a form of data splitting. Please, note that in this situation, the test data is only independent to the model building data in terms of all the data processing operations that we make AFTER data splitting, but it is not independent in terms of all the previous operations, included the ones involved in the data generation process. This is a limitation that should be recognized.
@@
There is a large number of criteria for performance evaluation [17], to give some few examples: Predictive Error Sum-of-Squares (PRESS), the Goodness of Prediction ($Q^{2}$), Mean Absolute Error (MAE), Mean Square Error (MSE), Precision, Recall, F1 Score, Area Under the Receiver Operating Characteristic (AUROC), Number of Misclassification (NMC), Mathews Coefficient Correlation (MCC), Cohen’s Kappa, etc. Quite often, when comparing multiple model variants, these criteria provide conflicting results. In real-life applications, however, there are considerations to make, often of practical nature, about when a solution is preferred over other solutions. For instance, there may be different perceived degrees of severity when making false positives and false negatives. In the diagnosis of a severe disease, a false positive may be an unpleasant experience for the patient, but a false negative may lead to very negative consequences for their health. In a critical infrastructure, a false positive of malfunctioning may lead to wasting the time of a technician, but a false negative may lead to the disruption of the critical service. Even two different types of false positives (or negatives) may not be perceived the same in a real application. Understanding the severity of inference errors in the real-life application of a model is extremely important for proper validation. The application experts should provide this information.
@@
Take the example of Figure 1, where a random $\mathbf{y}~(1000\times 1)$ is estimated by a hypothetical classifier $\hat{\mathbf{y}}~(1000\times 1)$ with some level of noise. We repeat the generation of $\mathbf{y}$ and $\hat{\mathbf{y}}$ ten times, emulating what would happen if we assess the performance of the classifier over ten different test datasets with the same imbalance level. When $\mathbf{y}$ is composed of a majority class with $70\%$ of the observations and a minority class with $30\%$, the resulting ROC curve is very stable (Figure 1(a), with an AUROC [^3] The Area Under the ROC curve or AUROC ranges from 0.5 for a random classifier to an optimal value of 1 for a perfect classifier. between 0.84 and 0.86). When we assess the same classifier with a minority class of $1\%$, which is realistic or even large for many anomaly detection problems or in the diagnosis of many diseases, the random variability associated to the ROC curve grows severely (Figure 1(b), with an AUROC between 0.77 and 0.89). Clearly, a ROC (or AUROC) is an unstable performance measure for this type of problems.
@@
Figure 1: Receiver Operating Characteristic (ROC) curves for unbalanced data. A ROC curve is a graphical plot that illustrates the diagnostic ability of a binary classifier system as its discrimination threshold is varied, created by plotting the True Positive Rate (Sensitivity) against the False Positive Rate (1 - Specificity) at various threshold settings.
@@
A different angle is considered in Figure 3. Rather than the NMC, we consider a weighted performance function where false positives are pondered with a unit and false negatives with 100 units, incorporating the idea that a false negative is something we would rather avoid at the expense of making more false positives. Back in the case with a minority class of $1\%$, this approach shows a better performance of the original classifier (in Figure 3(a)) than the naive approach that always outputs the negative class (in Figure 3(b)). This perceived performance is more consistent for a real-case problem where false negative are more relevant. Furthermore, this performance function shows that reducing the threshold for the detection (which generates less false negatives at the expense of generating more false positives) can further improve performance (Figure 3(c)). Clearly, the nature of the performance function can change both our perception of performance and the consequent choices we make to configure the optimal model. For this reason, choosing a performance criterion that is consistent with the real-life application is key to make adequate choices and have realistic assessment of the model quality.
@@ 2.4 Rule 4: Relevant baselines should always be provided
It is also a strong recommendation to report relevant performance baselines. A baseline can be selected from different perspectives. To give an example, popular cross-validation curves for the selection of the number of latent variables (LVs) in component models (like PCA or Partial Least Squares (PLS)) are too often shown without a baseline. In this case, a suggested baseline [18] can be the cross-validated performance criterion (e.g., the sum of squares) of the model with 0 LVs, that is, only considering the preprocessing parameters, like for instance the cross-validated means. Quite often, not a single model variant, regardless the number of LVs, yields an improvement over the 0 LVs baseline. This is a situation in which the selected component model is not an adequate choice, or there is simply no pattern to be found in the data. This situation cannot be detected if a baseline is not provided.
Take the example of Figure 4, with the cross-validation profile of a PLS model where $\mathbf{X}~(20\times 10)$ and $\mathbf{y}~(20\times 1)$ are unrelated. The cross-validation curves are based on the PRESS, defined as:
$PRESS=\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (1)
where $y_{i}$ represents each observation in $\mathbf{y}$ and $\hat{y}_{i}$ its corresponding prediction. In cross-validation, $\hat{y}_{i}$ is computed with a model which was not trained with $y_{i}$.
Figure 4(a) shows a cross-validation curve with a minimum at two LVs. An obvious baseline we are missing in this first plot is to fix the minimum of the Y scale at PRESS = 0. When we add it in Figure 4(b) we realize that the difference in PRESS between all the models, especially the ones with one LV and two LVs, is very small. Adding the zero LVs baseline, represented by the error of estimating $\mathbf{y}$ with its own (cross-validated) average, also provides useful information. This is illustrated in Figure 4(c), where we also show with dashed lines two hypothetical values for the PRESS at 0 LVs that would dramatically change our interpretation. If the PRESS at zero LVs is 100, the interpretation would be that the model with one or two LVs is actually capturing more than 60% of the variance in $\mathbf{y}$ in a predictive way. If the PRESS at zero LVs is 10, the interpretation would be that no PLS model has any predictive power. This is actually what the real value of the PRESS at zero LVs (around 30) is implying.
@@
Another form of baseline is represented by null examples. A null example is a negative case, fabricated using random number generators while maintaining as much as possible the characteristics of the real data at hand—like the number of rows and columns, group imbalances, selected correlation structures, etc. The null distribution in permutation testing [19], or the null effect size in powercurves [20], for instance, follow this intuition. The resulting performance reflects a baseline, which is often (much) higher than expected [21]. Assessing any data pipeline with one or several null examples is an excellent supplement for validation. Null examples are a perfect tool to detect data leakage problems, because they lead to an unexpectedly high performance when such problems exist.
Take now the example of Figure 5, similar as the previous example but for $\mathbf{X}~(20\times 1000)$ and $\mathbf{y}~(20\times 1)$. Since both blocks of data are unrelated, this represent a null example of a case where $\mathbf{X}$ is highly dimensional. A correct validation scheme should reflect the null predictive power of $\mathbf{X}$ over $\mathbf{y}$. Figure 5(a) shows the (incorrect) case where we do variable selection with the whole dataset and then perform double cross-validation. Figure 5(b) shows the (correct) case where we do variable selection within the inner loop of the double cross-validation. In both cases, we present one example of the cross-validation curve in the inner loop, and the final Goodness of prediction ($Q^{2}$) of the selected model, defined as:
$Q^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{SS_{y}}$ (2)
with $SS_{y}$ the sum-of-squares of $\mathbf{y}$. In double cross-validation, $\hat{y}_{i}$ is computed with a model which was neither trained nor selected with $y_{i}$.
The null example shows the overly optimistic nature of the incorrect validation scheme in Figure 5(a), where the optimal model has one LV and a fairly good (perceived) prediction performance. The correct validation scheme in Figure 5(b) shows that $\mathbf{X}$ does not contain information on $\mathbf{y}$. The nice property of a null example is that it is extremely easy to generate, either from random data or by permutation of a real dataset, but very powerful to detect flawed validation schemes. In very complex pipelines, with several steps for preprocessing, feature engineering and modeling, running some permuted examples is an effortless way to double-check that the pipeline is not suffering from any data leakage problem.
@@
Figure 5: Cross-validation curve and Goodness of prediction ($Q^{2}$) of double cross-validation for Partial Least Squares (PLS) in a simulated dataset where $\mathbf{X}(20\times 1000)$ and $\mathbf{y}(20\times 1)$ are unrelated: variable selection performed before validation (a) and variable selection performed within the inner loop (b).